OpenStudy (anonymous):

Financial Math Help Stan makes a single deposit of $5000 into an account that earns 6% compounded quarterly. How much will Stan have at the end of 10 years?

OpenStudy (koalagurl1):

$12,000

OpenStudy (koalagurl1):

%6 of $5000 is $300

OpenStudy (koalagurl1):

$300*4=$1200

OpenStudy (koalagurl1):

$1200*10=$12000

OpenStudy (anonymous):

well yeah that sounds right but im supposed to use an certain equation (work is required to get credit in our homework assignments) Ill just use trial and error until i come out with 12000 as my answer i guess haha

OpenStudy (wolf1728):

Sorry KoalaGurl1 - here's how it's done. First of all we have to convert 6% compounded quarterly to an annual rate. Annual Rate = (1+ .06/4)^4 Annual Rate =1.0613635506 Now we can go to the compound interest formula: Total = Principal*(1 + rate)^years

OpenStudy (anonymous):

i actually think the answer is $8163.08 equation i think im supposed to use is \[M=P(1+\frac{ R }{ N })^{(NT)}\] with P=5000 R/N = .015 (NT) = 40

OpenStudy (anonymous):

what you think wolf? does that look right?

OpenStudy (wolf1728):

Total = Principal*(1 + rate)^years Total = 5,000 * (1.0613635506)^10 =9,070.0920433447 which rounds to $9,070.09

OpenStudy (anonymous):

awh man, so i was wrong

OpenStudy (wolf1728):

Okay nseller23 , I know this stuff. Need any explanation?

OpenStudy (anonymous):

oh wait, i was using wrong numbers. i did my actual homework problem, i posted a similar problem with different values lol. now this makes more sense

OpenStudy (anonymous):

lemme check with the values i posted.

OpenStudy (koalagurl1):

:(

OpenStudy (anonymous):

alright yes wolf, i got same answer. so we did same way

OpenStudy (anonymous):

Ive got 10 different problems for my homework assignment, i may tag your name if i cant get them as i move along.

OpenStudy (koalagurl1):

sorry, I'm in 7th grade and theres not a lot I can answer on here

OpenStudy (wolf1728):

Okay then!

OpenStudy (wolf1728):

That's okay KoalaGurl1 - you tried

OpenStudy (anonymous):

no koala i probably would have gone after the same method as you if i wasnt taking the class haha. good effort and keep that up

OpenStudy (koalagurl1):

thx :)

OpenStudy (koalagurl1):

It was too good to be true lol

OpenStudy (anonymous):

haha heard that :D

OpenStudy (koalagurl1):

:)

OpenStudy (wolf1728):

By the way nsellers23, here's a compound interest calculator to check your work: http://www.1728.org/compint.htm

OpenStudy (anonymous):

oh wow thank you! haha youre a life saver

OpenStudy (wolf1728):

thanks nseller23

OpenStudy (koalagurl1):

I like your pfp btw :P @wolf1728

OpenStudy (wolf1728):

What is pfp?

OpenStudy (koalagurl1):

profile picture

OpenStudy (wolf1728):

Oh wow - thanks! A lot of people tell me that! (I like wolves).

OpenStudy (koalagurl1):

yeah me too

OpenStudy (wolf1728):

Thanks !

OpenStudy (koalagurl1):

:) np

OpenStudy (anonymous):

wolf i have a very similar problem but got me confused. Stan deposits $4500 at the end of EACH quarter at 6% compounded quarterly. How much will hakeem have after 10 yrs? Did i get the correct answer: $244,205.52

OpenStudy (anonymous):

wait no, thats wrong.

OpenStudy (anonymous):

maybe its right, idk. i realized i made a mistake but came up with the same answer. please check if you dont mind.

OpenStudy (wolf1728):

Okay, that is considered an annuity. (I'm working on this).

OpenStudy (anonymous):

would it be an ordinary annuity?

OpenStudy (wolf1728):

I would think so

OpenStudy (anonymous):

yeah thats what im thinking aswell. seeing a similar example under the ordinary annuity section.

OpenStudy (anonymous):

btw do you need any help with anything or you just on here helping others?

OpenStudy (wolf1728):

I don't need help. (I only help 1 person at a time).

OpenStudy (anonymous):

awesome. i like that. If its the same problem but deposited at the beginning of each quarter, does that change anything?

OpenStudy (wolf1728):

I suppose it would change things if it were for the beginning of each quarter. As of the moment I'm working on determining a quarterly annuity formula. Do you have one?

OpenStudy (anonymous):

the formula i have for an ordinary annuity is